OpenStudy (anonymous):

how to do I x+4 I <= I x - 1 I ?

OpenStudy (amistre64):

might help to find out when they are equal to start with, then we can deduce it further

OpenStudy (anonymous):

im pretty sure we have to get like 3 possible solutions or something like that.

OpenStudy (amistre64):

x+4 = x-1 x+4 = -(x-1) these equivalences should tell us when they are equal

OpenStudy (anonymous):

i know that you have to do something like....